OpenStudy (anonymous):

Whats the radius of the equation: (x-2)^2 + y^2 = 14

OpenStudy (ranga):

The general equation of a circle (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center of the circle and r is the radius. Compare (x-2)^2 + y^2 = 14 to the above and you can quickly find the radius r.

OpenStudy (anonymous):

I dont know how to convert it into general form.

OpenStudy (ranga):

You don't have to convert it. They have already given you the equation in general form. You simply have to compare them: (x-h)^2 + (y-k)^2 = r^2 (x-2)^2 + (y-0)^2 = 14 You can see that r^2 = 14. Therefore the radius r is sqrt(14).